OpenStudy (anonymous):

A part time employee worked 4 hours on monday and 7 hours each day for the next 3 days. Write and simplfy an expression that the total numbers of hours worked

OpenStudy (gw2011):

Total=4+(7)(3)=4+21=25 hours worked